Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Twenty people at a meeting were born during the month of

Expert replies
by AAPL » Tue Jan 15, 2019 4:57 am
Timer
00:00
Your answer
Difficulty–

right wrong

Pick an answer to see how members did

Veritas Prep

Twenty people at a meeting were born during the month of September, which has 30 days. The probability that at least two of the people in the room share the same birthday is closest to which of the following?

A. 10%
B. 33%
C. 67%
D. 90%
E. 99%

OA E
Join the discussion
Source: — Problem Solving |

by fskilnik@GMATH » Tue Jan 15, 2019 5:29 am
AAPL wrote:Veritas Prep

Twenty people at a meeting were born during the month of September, which has 30 days. The probability that at least two of the people in the room share the same birthday is closest to which of the following?

A. 10%
B. 33%
C. 67%
D. 90%
E. 99%
$$? = 1 - P\left( {\underbrace {{\rm{all}}\,\,20\,\,{\rm{different}}\,\,{\rm{birthday}}\,\,{\rm{dates}}}_{{\rm{unfavorable}}}} \right)$$

$${\rm{Total}}:\,\,30 \cdot 30 \cdot \ldots \cdot 30 = {30^{20}}\,\,\,{\rm{equiprobable}}\,\,{\rm{possibilities}}\,\,\,$$
$${\rm{unfavorable}} = \,\,30 \cdot 29 \cdot \ldots \cdot 11$$
$$P\left( {{\rm{unfavorable}}} \right) = {{30 \cdot 29 \cdot \ldots \cdot 11} \over {{{30}^{20}}}} = 1 \cdot \underbrace {{{29} \over {30}} \cdot {{28} \over {30}} \cdot \ldots {{14} \over {30}}}_{ < < < \,\,1} \cdot \underbrace {{{13} \over {30}} \cdot {{12} \over {30}} \cdot {{11} \over {30}}}_{ \cong \,\,0.05} < < < 0.05 = 5\% $$
$$?\,\,\, > > > \,\,\,100\% - 5\% \,\,\,\,\,\,\, \Rightarrow \,\,\,\,\,\,\left( {\rm{E}} \right)$$


This solution follows the notations and rationale taught in the GMATH method.

Regards,
Fabio.
Fabio Skilnik :: GMATH method creator ( Math for the GMAT)
English-speakers :: https://www.gmath.net
Portuguese-speakers :: https://www.gmath.com.br
Join the discussion

by Jay@ManhattanReview » Tue Jan 15, 2019 5:45 am
AAPL wrote:Veritas Prep

Twenty people at a meeting were born during the month of September, which has 30 days. The probability that at least two of the people in the room share the same birthday is closest to which of the following?

A. 10%
B. 33%
C. 67%
D. 90%
E. 99%

OA E
Probability that at least two people sharing the same birthday = 1 – [Probability that NONE of them sharing the same birthday]

Number of ways of NONE of the 20 shares the birthday = 30P20 = 30!/(30–20)! = 30!/10! = 11*12*......*29*30; we chose permutation since the order of selection matters.

The total number of possible ways of 20 people born in September = 30*30*......*30*30 = 30^20 (Each of the 20 has 30 options);

Probability that at least two people sharing the same birthday = 1 – [(11*12*......*29*30)/(30^20)]

Note that in (11*12*......*29*30)/(30^20), the numerator is significantly less than the denominator; thus, (11*12*......*29*30)/(30^20) is very less.

Or, Probability that at least two people sharing the same birthday = 1 – [(11*12*......*29*30)/(30^20)] = ~0.99.

The correct answer: E

Hope this helps!

-Jay
_________________
Manhattan Review

Locations: Manhattan Review Jayanagar | GMAT Prep Tarnaka | GRE Prep Madhapur | Kukatpally GRE Coaching | and many more...

Schedule your free consultation with an experienced GMAT Prep Advisor! Click here.
Join the discussion

by Scott@TargetTestPrep » Thu Jan 17, 2019 5:58 pm
AAPL wrote:Veritas Prep

Twenty people at a meeting were born during the month of September, which has 30 days. The probability that at least two of the people in the room share the same birthday is closest to which of the following?

A. 10%
B. 33%
C. 67%
D. 90%
E. 99%
The probability that at least two of the people in the room share the same birthday is equivalent to subtracting from 1 the probability that no two people in the room share the same birthday.

The first person can have a birthday on any of the 30 days of September. In order to avoid a birthday match, the second person can have a birthday on any of the remaining 29 days. Similarly, to avoid a match with either of the first two people, the third person can have a birthday on any of the remaining 28 days. And so forth, down to the twentieth person. We can then express each event as a probability by dividing by 30, the total number of days in September. The first person’s probability of not matching is 30/30 (because they can be born on any day). The second person’s probability of not matching the first person is 29/30, and the third person’s probability of not matching either of the first two is 28/30. This follows in a similar fashion to the twentieth person.

The probability that no two people in the room share the same birthday (i.e., that they all have different birthdays) is:

30/30 x 29/30 x 28/30 x … x 11/30
(30 x 29 x 28 x … x 11)/(30 x 30 x 30 x … x 30)

30P20 / 30^20 ≈ 0.0002

Therefore, the probability that at least two of the people in the room do share the same birthday is:

1 - 0.0002 = 0.9998 = 99.98%

Answer: E

Scott Woodbury-Stewart
Founder and CEO
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews

ImageImage
Join the discussion